# Center for Advanced Intelligence Project, RIKEN 103-0027, Tokyo, Japan

### v1

@@ 4.2 Univariate representations of the solutions
Let $f$ be $\operatorname{\mathcal{D}}_{0}$. Then, there exist polynomials $w,v_{1},\dots,v_{n}$ in $\operatorname{\mathbb{R}}[t]$, with $\deg v_{i}<\deg w$, $i\in[n]$, and $\deg w=\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$, such that
@@
Since $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])$ is finite, according to (rouillier1999, , Lemma 2.1), there exists an invertible $3n\times 3n$-matrix $H$ with real entries such that, if we change the variables $(x,z)=Hy$ into the following polynomials
@@
obtain respectively polynomials $h_{1}(y),\dots,h_{3n}(y)$, and denote by the ideal in $\operatorname{\mathbb{R}}[y_{1},\dots,y_{3n}]$
@@
Clearly, the ideal $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ satisfies the condition of the Shape Lemma. So, there exist $3n$ polynomials in $\operatorname{\mathbb{R}}[y_{1}]$, says $w$, $u_{2},\dots,u_{3n}$, with $\deg w=\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$ and $\deg u_{i}<\deg w$, $i=2,\dots,3n$, and such that $[w,y_{2}-u_{2},\dots,y_{3n}-u_{3n}]$ is the reduced Gröbner basis with respect to the lexicographic order with $y_{1}<\cdots<y_{3n}$ of $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$.
@@
Recall that the entries of $H$ are real, then that ones of $H^{-1}$ so are. Suppose that $\tau$ is a root of $w(t)$. If $\tau$ is real, then all coordinates of $H(u(\tau))^{T}$ are real, hence this vector belongs to $V_{\operatorname{\mathbb{R}}}(\operatorname{\mathcal{I}}[f,z])$. Conversely, if $H(u(\tau))^{T}$ in $\operatorname{\mathbb{R}}^{3n}$ then
@@
Let $v_{1}(t),\dots,v_{n}(t)$ be the first $n$ coordinates of $H(u(t))$. Since polynomials $u_{2},\dots,u_{3n}$ have real coefficients, so do $v_{1},\dots,v_{n}$. Moreover, if $\deg v_{i}\geq\deg w$ then we replace $v_{i}$ by the remainder of the division of $v_{i}$ by $w$, for $i\in[n]$.
@@
The system $[w,v_{1},\dots,v_{n}]$ mentioned in Theorem 4.1 is univariate and such a system is called to as a univariate representation of the solutions of $\operatorname{CP}(f)$.
@@
we obtain six polynomials $h_{1},\dots,h_{6}$ in variables $y$. The ideal $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ is in shape position with the graded lexicographic order with $y_{1}<y_{2}<\cdots<y_{6}$, with the Gröbner basis of $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$, say $[w(y_{1}),y_{2}-u_{2}(y_{1}),\dots,y_{6}-u_{6}(y_{1})]$. The first two coordinates of $H(t,u_{2},\dots,u_{6})$ are denoted by $v_{1},v_{2}$. So we obtain $w$ and $v_{1},v_{2}$ as follows:
$w(t)=t^{8}+4t^{7}+78t^{6}+392t^{5}-5247t^{4}-11228t^{3}-5324t^{2}-11616t$, $v_{1}(t)=(-395708t^{7}-1758766t^{6}-31180792t^{5}-167371327t^{4}+2037780580t^{3}+5545360564t^{2}+2068565664t+5710973125)/5710973125$, $v_{2}(t)=(395708t^{7}+1758766t^{6}+31180792t^{5}+167371327t^{4}-2037780580t^{3}-5545360564t^{2}-2068565664t-5710973125)/5710973125$.
Polynomial $w$ has eight roots including $\{-8,-2,0,6,i,-i,11i,-11i\}$. According to (4.3), the solution set is given by
$\operatorname{Sol}(f)=\big{\{}(v_{1}(t),v_{2}(t)),t\in\{-8,-2,0,6\}\big{\}}=\{(1,1)\}.$
@@ 4.3 Univariate representations of least-norm solutions and sparse solutions
is a univariate polynomial with real coefficients. Denote by $\varphi$ the remainder of the division of $\hat{\varphi}$ by $w$. One has $\deg\varphi<\deg w$. Let $\gamma$ be the minimum of $\varphi(t)$ over the finite set $\{t\in\operatorname{\mathbb{R}},w(t)=0\}$. It is not difficult to prove that $\hat{\alpha}$ is a least-norm solution of $\operatorname{CP}(f)$ if and only if there exists a real root $\tau$ of $w$ such that $\varphi(\tau)=\gamma$ and $\hat{\alpha}=(v_{1}^{2}(\tau),\dots,v_{n}^{2}(\tau))$. Hence, (4.6) is proved. ∎
@@
Let $[w,v_{1},\dots,v_{n}]$ be a univariate representation of $\operatorname{Sol}(f)$ and $\alpha$ be a solution of $\operatorname{CP}(f)$. Let $\ell$ be a subset of $[n]$. We denote by ¯$\bar{\ell}=[n]\setminus\ell$ and $w_{\ell}$ by the greatest common divisor defined as follows:
@@
$\operatorname{Sol_{\operatorname{sp}}}(f)=\bigcup_{\ell\subseteq[n],k=\#\ell}\Big{\{}\big{(}v_{1}(t),\dots,v_{n}(t)\big{)}:w_{\ell}(t)=0,w_{\bar{\ell}}(t)\neq 0,t\in\operatorname{\mathbb{R}}\Big{\}}.$ (4.8)
@@ 5.1 Computing univariate representations of $\operatorname{Sol}(f)$
Input and output. The input of Algorithm 1 is a polynomial mapping $f$ being $\operatorname{\mathcal{D}}_{0}$. The output includes a list of univariate polynomials $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ such that $\operatorname{Sol}(f)$ can be represented through the univariate polynomials as in (5.1).
@@
Output: $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ a univariate representation of $\operatorname{Sol}(f)$
@@
      Change variables $(x,z)=Hy$ in the polynomials in (4.4), respectively, obtain polynomials denoted by $g_{i}(y)$, $i\in[3n]$, where
@@
      Set $\operatorname{\mathcal{I}}_{h}\coloneqq\left\langle h_{1},\dots,h_{3n}\right\rangle$ in $\operatorname{\mathbb{R}}[y_{1},\dots,y_{3n}]$ and compute the reduced Gröbner basis $G$ of $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ with respect to the lexicographic order $y_{1}<y_{2}<\cdots<y_{3n}$
@@
      $G=[\hat{w},y_{2}-u_{2},\dots,y_{3n}-u_{3n}],$ (5.3)
      where $\hat{w},u_{2},\dots,u_{3n}$ are in $\operatorname{\mathbb{R}}[y_{1}]$, i.e. $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ is in shape position, then set $\operatorname{\mathtt{shape}}\coloneqq$ true, else set $s\coloneqq s+1$
@@
    Define by $(v_{1},\dots,v_{n})$ the first $n$ coordinates of $H(t,u_{2},\dots,u_{3n})$
@@
Description. At the beginning, we set $s=0$ and $\operatorname{\mathtt{shape}}=$ false. The variable $\operatorname{\mathtt{shape}}$ is a boolean. The while loop at Line 2 will stop if $\operatorname{\mathtt{shape}}$ is true. Within the while loop, at Line 3, with $s$ given, we define matrix $H$ depending on $s$ as in (5.2) and change the variables $(x,z)=Hy$ in the polynomials in (4.4) and obtain $h_{i}$, $i\in[3n]$. At Line 4, we set
@@
and compute the reduced Gröbner basis $G$ of the radical of the ideal $\operatorname{\mathcal{I}}_{h}$, that is $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$, with respect to the lexicographic order with $y_{1}<y_{2}<\cdots<y_{3n}$. At Line 5, we set the value of $\operatorname{\mathtt{shape}}$ as true if $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ is in shape position otherwise, we increase the value $s\coloneqq s+1$.
@@
Denote by $\delta$ the cardinality of $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])$, according to (rouillier1999, , Lemma 2.1), there is a number $j$ in the set $\{0,\dots,(3n-1)\delta(\delta-1)/2\}$ such that the linear function $u\coloneqq\left\langle(x,z)\mid(1,s,\dots,s^{3n-1})\right\rangle$ separates $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$, i.e., $u(a)\neq u(b)$ for any distinct points $a,b$ in $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$. After changing the variables at Line 3, namely, $(x,z)=Hy$, the radical of $\operatorname{\mathcal{I}}_{h}$ satisfies the assumptions of Lemma 2. Ideal $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ is in shape position and the while loop ends within at most $(3n-1)\delta(\delta-1)/2+1$ iterations.
After the while loop, the reduced Gröbner basis of $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ written as at Line 5, $G=[w,y_{2}-u_{2},\dots,y_{n}-u_{3n}]$. After Line 6, one has
@@
From the latter equality and (1), we get (5.1). Therefore, we claim that the output of Algorithm 1 provides a univariate representation of $\operatorname{CP}(f)$. ∎
@@
However, in practice the matrix $H$ should be chosen randomly with rational entries, hence we do not need to use the while loop at Line 2.
@@ 5.2 Computing univariate representations of $\operatorname{Sol_{\operatorname{ln}}}(f)$
Input: $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ a univariate representation of $\operatorname{Sol}(f)$, where $f$ is $\operatorname{\mathcal{D}}_{0}$
@@ 5.3 Computing univariate representations of $\operatorname{Sol_{\operatorname{sp}}}(f)$
We design Algorithm 3 to compute a univariate representation of the sparse solution set of a $\operatorname{\mathcal{D}}_{0}$ problem. The output of Algorithm 3 consists of a subset $\Omega\subset 2^{[n]}$ and $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ such that, for every element $\ell$ in $\Omega$, the following set is nonempty:
@@
Input: $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ a univariate representation of $\operatorname{Sol}(f)$, where $f$ is $\operatorname{\mathcal{D}}_{0}$
@@
Description. In Algorithm 3, $k$ indicates the number of zero coordinates of a sparse solution, and $\Omega\subset 2^{[n]}$ is displayed as the set of all $\ell\subset[n]$ with #$\#\ell=k$ such that there is a sparse solution whose coordinates at $\ell$ are zero. In the initial stage, we set $k=0$ and $\Omega=\emptyset$, which indicates that the problem lacks sparse solutions.
Since $k$ belongs to $\{0,1,\dots,n\}$ and is the maximum number of zero coordinates, in Line 2 we reduce the value of $j$ starting from $n$. The for loop is stopped when the first index set $\ell$, of a sparse solution, is found. We note that, for each $j$, there are $n!/(j!(n-j)!)$ subset of $[n]$ with cardinality $j$. Once a sparse solution is found, in line 3, for each $\ell\subseteq[n]$ with $k$ elements, if $w_{\ell}$ has real roots and $w_{\bar{\ell}}$ has no real roots, then we let the index set to belong to $\Omega$.
Correctness. The correctness of Algorithm 3 is briefly explained as follows: After the for loop in Line 2, $k$ is the value of $\|\alpha\|_{0}$ for a sparse solution $\alpha$. Clearly, $\ell\in\Omega$ if and only if the set (5.4) is non-empty. Furthermore for each $\tau$ in the set (5.4), $(v_{1}(\tau),\dots,v_{n}(\tau))$ is a sparse solution whose coordinates at $\ell$ are zero. Hence, we get the equality (5.5).

### v3

@@ 4.3 Univariate Representations of the Solutions
Let $f$ be $\operatorname{\mathcal{D}}_{0}$. Then, there exist univariate polynomials $v_{1},\dots,v_{n}$, $w$ in $\operatorname{\mathbb{R}}[t]$, with $\deg v_{i}<\deg w$, $i\in[n]$, and $\deg w=\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$, such that
@@
Since $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])$ is finite, according to (rouillier1999, , Lemma 2.1), there exists an invertible $3n\times 3n$-matrix $H$ with real entries such that, if we change the variables by $(x,z)=Hy$ in the following polynomials
@@
we obtain respectively polynomials $h_{1}(y),\dots,h_{3n}(y)$, and introduce in $\operatorname{\mathbb{R}}[y_{1},\dots,y_{3n}]$ the ideal
@@
Clearly, the ideal $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ satisfies the condition of the Shape Lemma. So, there exist $3n$ polynomials in $\operatorname{\mathbb{R}}[y_{1}]$, say $w$, $u_{2},\dots,u_{3n}$, with $\deg w=\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$ and $\deg u_{i}<\deg w$, $i=2,\dots,3n$, and such that $[w,y_{2}-u_{2},\dots,y_{3n}-u_{3n}]$ is the reduced Gröbner basis of $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ with respect to a lexicographic monomial ordering $y_{1}<\cdots<y_{3n}$.
@@
Recall that the entries of $H$ are real, so that the ones of $H^{-1}$ are so as well. Suppose that $\tau$ is a root of $w(t)$. If $\tau$ is real, then all coordinates of $H(u(\tau))^{T}$ are real, hence this vector belongs to $V_{\operatorname{\mathbb{R}}}(\operatorname{\mathcal{I}}[f,z])$. Conversely, if $H(u(\tau))^{T}$ is in $\operatorname{\mathbb{R}}^{3n}$ then
@@
Let $v_{1}(t),\dots,v_{n}(t)$ be the first $n$ components of $H(u(t))$. Since the polynomials $u_{2},\dots,u_{3n}$ have real coefficients, so do $v_{1},\dots,v_{n}$. Moreover, if $\deg v_{i}\geq\deg w$ then we replace $v_{i}$ by the remainder of the division of $v_{i}$ by $w$, for $i\in[n]$.
@@
The system $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ mentioned in Theorem 4.2 is univariate and such a system is called a univariate representation of the solutions of $\operatorname{CP}(f)$.
@@
we obtain six polynomials $h_{1},\dots,h_{6}$ in variables $y$. The ideal $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ is in shape position with the pure lexicographic ordering $y_{1}<y_{2}<\cdots<y_{6}$. The Gröbner basis of $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ is $[w(y_{1}),y_{2}-u_{2}(y_{1}),\dots,y_{6}-u_{6}(y_{1})]$. The first two coordinates of $H(t,u_{2},\dots,u_{6})$ are denoted by $v_{1},v_{2}$. So we obtain $w$ and $v_{1},v_{2}$ as follows:
$w(t)=t^{8}+4t^{7}+78t^{6}+392t^{5}-5247t^{4}-11228t^{3}-5324t^{2}-11616t$, $v_{1}(t)=(-395708t^{7}-1758766t^{6}-31180792t^{5}-167371327t^{4}+2037780580t^{3}+5545360564t^{2}+2068565664t+5710973125)/5710973125$, $v_{2}(t)=v_{1}(t)$.
Polynomial $w$ has eight roots $-8,-2,0,6,\mathrm{i},-\mathrm{i},11\mathrm{i},-11\mathrm{i}$, where $\mathrm{i}$ is the imaginary unit. According to (4.3), the solution set is given by
$\operatorname{Sol}(f)=\big{\{}(v_{1}(t),v_{2}(t)):t\in\{-8,-2,0,6\}\big{\}}=\{(1,1)\}.$
@@ 4.4 Univariate Representations of Least-norm Solutions and Sparse Solutions
is a univariate polynomial with real coefficients. Denote by $\varphi$ the remainder of the division of $\hat{\varphi}$ by $w$. One has $\deg\varphi<\deg w$. Let $\gamma$ be the minimum of $\varphi(t)$ over the finite set $\{t\in\operatorname{\mathbb{R}},w(t)=0\}$. It is not difficult to prove that $\hat{\alpha}$ is a least-norm solution of $\operatorname{CP}(f)$ if and only if there exists a real root $\tau$ of $w$ such that $\varphi(\tau)=\gamma$ and $\hat{\alpha}=(v_{1}(\tau),\dots,v_{n}(\tau))$. Hence, (4.6) is proved. ∎
@@
Let $[w,v_{1},\dots,v_{n}]$ be a univariate representation of $\operatorname{Sol}(f)$ and $\alpha$ be a solution of $\operatorname{CP}(f)$. Let $\ell$ be a subset of $[n]$. We denoted $w_{\ell}$ by the greatest common divisor defined as follows:
@@
$\operatorname{Sol_{\operatorname{sp}}}(f)=\bigcup_{\ell\subseteq[n],k=l}\Big{\{}\big{(}v_{1}(t),\dots,v_{n}(t)\big{)}:w_{\ell}(t)=0,w_{\bar{\ell}}(t)\neq 0,t\in\operatorname{\mathbb{R}}\Big{\}}.$ (4.8)
@@ 5.1 Computing Univariate Representations of $\operatorname{Sol}(f)$
Input and output. The input of Algorithm 1 is a polynomial mapping $f$ being $\operatorname{\mathcal{D}}_{0}$. The output is a list of univariate polynomials $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ such that $\operatorname{Sol}(f)$ can be represented as in (5.1).
@@
Output: $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ a univariate representation of $\operatorname{Sol}(f)$,
@@
      Change variables $(x,z)=Hy$ in the polynomials in (4.4), respectively, obtain polynomials denoted by $h_{i}(y)$, $i\in[3n]$, where
@@
      Set $\operatorname{\mathcal{I}}_{h}\coloneqq\left\langle h_{1},\dots,h_{3n}\right\rangle$ in $\operatorname{\mathbb{R}}[y_{1},\dots,y_{3n}]$ and compute the reduced Gröbner basis $G$ of $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ with respect to the pure lexicographic ordering $y_{1}<y_{2}<\cdots<y_{3n}$
@@
      $G=[w,y_{2}-u_{2},\dots,y_{3n}-u_{3n}],$ (5.3)
      where $w,u_{2},\dots,u_{3n}$ are in $\operatorname{\mathbb{R}}[y_{1}]$, i.e. $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ is in shape position, then set $\operatorname{\mathtt{shape}}\coloneqq$ true, else set $s\coloneqq s+1$
@@
    Change the variable $y_{1}$ by $t$ and define by $(v_{1},\dots,v_{n})$ the first $n$ coordinates of $H(t,u_{2},\dots,u_{3n})^{T}$
@@
Description. At the beginning, we set $s=0$ and $\operatorname{\mathtt{shape}}=$ false. The variable $\operatorname{\mathtt{shape}}$ is boolean. The while loop in Line 2 will stop if $\operatorname{\mathtt{shape}}$ is true. Within the while loop, in Line 3, with $s$ given, we define matrix $H$ depending on $s$ as in (5.2) and change the variables $(x,z)=Hy$ in the polynomials in (4.4) and obtain $h_{i}$, $i\in[3n]$. In Line 4, we set
@@
and compute the reduced Gröbner basis $G$ of the radical of the ideal $\operatorname{\mathcal{I}}_{h}$, that is $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$, with respect to the pure lexicographic order with $y_{1}<y_{2}<\cdots<y_{3n}$. In Line 5, we set the value of $\operatorname{\mathtt{shape}}$ as true if $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ is in shape position otherwise, we increase the value $s\coloneqq s+1$.
@@
Denote by $\delta$ the cardinality of $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])$, according to (rouillier1999, , Lemma 2.1), there is a number $s$ in the set $\{0,\dots,(3n-1)\delta(\delta-1)/2\}$ such that the linear function
$u(x,z)\coloneqq x_{1}+\dots+s^{n-1}x_{n}+s^{n}z_{1}+\dots+s^{3n-1}z_{2n}$
separates $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$, i.e., $u(a)\neq u(b)$ for any distinct points $a,b$ in $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$. Hence, the while loop ends within at most $(3n-1)\delta(\delta-1)/2+1$ iterations.
After the while loop in Line 2, the reduced Gröbner basis of $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$, written as in Line 5, is $G=[w,y_{2}-u_{2},\dots,y_{n}-u_{3n}]$. After Line 6, one has
@@
From the latter equality and (3.4), we get (5.1). Therefore, we claim that the output of Algorithm 1 is a univariate representation of $\operatorname{CP}(f)$. ∎
@@
However, in practice the matrix $H$ should be chosen randomly with rational entries, hence we do not need to use the while loop in Line 2.
@@ 5.2 Computing Univariate Representations of $\operatorname{Sol_{\operatorname{ln}}}(f)$
    Compute $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ a univariate representation of $\operatorname{Sol}(f)$
  • 2:
@@ 5.3 Computing Univariate Representations of $\operatorname{Sol_{\operatorname{sp}}}(f)$
We design Algorithm 3 to compute a univariate representation of the sparse solution set of a $\operatorname{\mathcal{D}}_{0}$ problem.
Description. The output of Algorithm 3 consists of a subset $\Omega\subset 2^{[n]}$ and $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ such that, for every element $\ell$ in $\Omega$, the following set is nonempty:
@@
    Compute $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ a univariate representation of $\operatorname{Sol}(f)$
  • 2:
@@
In Algorithm 3, $k$ indicates the number of zero coordinates of a sparse solution, and $\Omega\subset 2^{[n]}$ is displayed as the set of all $\ell\subset[n]$ with #$\#\ell=k$ such that there is a sparse solution whose coordinates at $\ell$ are zero. In the initial stage, we set $k=0$ and $\Omega=\emptyset$, which indicates that the problem lacks sparse solutions.
Since $k$ belongs to $\{0,1,\dots,n\}$ and is the maximum number of zero coordinates, in Line 3 we reduce the value of $j$ starting from $n$. The for loop is stopped when the first index set $\ell$, of a sparse solution, is found. We note that, for each $j$, there are $n!/(j!(n-j)!)$ subsets of $[n]$ with cardinality $j$. Once a sparse solution is found, in Line 4, we for each $\ell\subseteq[n]$ with $k$ elements, if $w_{\ell}$ has real roots and $w_{\bar{\ell}}$ has no real roots, add this set to $\Omega$.
Correctness. The correctness of Algorithm 3 is briefly explained as follows: After the for loop in Line 3, $k$ is the value of $\|\alpha\|_{0}$ for a sparse solution $\alpha$. Clearly, $\ell\in\Omega$ if and only if the set (5.4) is non-empty. Furthermore for each $\tau$ in the set (5.4), $(v_{1}(\tau),\dots,v_{n}(\tau))$ is a sparse solution whose coordinates at $\ell$ are zero. Hence, we get the equality (5.5).